Question 4 (8 points) – Based on the rule of 72 (your text page 21), how many years will it take for a $1,000 investment to grow to $2,000, assuming the investment grows at 6% per year?
Answer #1

According to the rule of 72;

n = 72/r = 72 / 6 = 12

So, it will take 12 years for $1,000 investment to grow to $2,000